OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 13, c = 21

OpenStudy (anonymous):

15.83 15.87 16.19 16.49

OpenStudy (cggurumanjunath):

c^2=a^2+b^2

OpenStudy (cggurumanjunath):

21^2 = a^2 +13^2

OpenStudy (cggurumanjunath):

21^2-13^2 =21*21-13*13 =? @Cookie_009

OpenStudy (anonymous):

wich one is it?

OpenStudy (cggurumanjunath):

21*21 -13*13 = ?

OpenStudy (anonymous):

272

OpenStudy (cggurumanjunath):

good

OpenStudy (cggurumanjunath):

a= sqrt (272)

OpenStudy (anonymous):

15.83 15.87 16.19 16.49

OpenStudy (anonymous):

its supose toi be one of thoseee

OpenStudy (cggurumanjunath):

USE A CALCULATOR.

OpenStudy (anonymous):

what?

OpenStudy (cggurumanjunath):

SQRT ( 272) = 16.49 BY USING A CALCULATOR WE GET THAT RESULT. THAT IS WHAT I MEANT. :)

OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 18, b = 26, c = ?

OpenStudy (cggurumanjunath):

proceed similarly .

OpenStudy (cggurumanjunath):

c^2 =a^2 + b^2